OpenStudy (anonymous):

Solve the equation x^2 + 5x = 0. a)x = 0, –5 b)x = 0, 5 c)x = 1, –5 d)x = 1, 5

OpenStudy (armyrangers):

Dose it look like \(\large x^2+5x\)

OpenStudy (anonymous):

test the first option if x =0, then 0^2+5*0 =0 check if x =-5, then (-5)^2 +5*(-5) = ??

OpenStudy (anonymous):

well there are many ways to solve this problem but an easy way would be to plug in solutions. For the first one: A) x=0,-5 x^2+5x=0 (0)^2+5(0)=0 0+0=0 0=0 good so far lets move on to -5 x^2+5x=0 (-5)^2+5(-5)=0 25-25=0 0=0 So A is the solution because both would work